OpenStudy (bossimbacon):

How many solutions are there to the following system of equations? 3x – 9y = 0 –x + 3y = –3 A. 0 B. 1 C. infinitely many D. 2

OpenStudy (bossimbacon):

@Directrix

OpenStudy (bossimbacon):

@DanJS

Directrix (directrix):

You can write the lines in the y = mx + b format. y = x/3 + 0 and y = x/3 - 1 and observe that the slopes are the same which means that the lines are parallel. The y-intercepts are different so the two lines are distinct. Or, you can observe that the lines are parallel here: http://www.wolframalpha.com/input/?i=3x+%E2%80%93+9y+%3D+0%2C+%E2%80%93x+%2B+3y+%3D+%E2%80%933

Directrix (directrix):

Of course, the question remains: How many solutions, if any, does the system have? @Bossimbacon Post what you think.

OpenStudy (bossimbacon):

@midhun.madhu1987

OpenStudy (midhun.madhu1987):

I agree with @Directrix If 2 lines are parallel, then slopes are equal. Here try to find the slopes of the 2 lines... by converting both equations to the form y=mx+b